# GaussianArt: Unified Modeling of Geometry and Motion for Articulated Objects

### v1

@@ 16.1 Experiments on PARIS Two-Part Dataset
Implementation Details. Following [65], we compare against Ditto [17], PARIS [33], PARIS* (PARIS augumented with depth supervision), CSG-reg [65], 3Dseg-reg [65], DigitalTwinArt [65], and ArtGS [38]. For a fair evaluation, we follow DigitalTwinArt and report the mean std for each metric over the 10 trials at the high-visibility state.
@@ 16.2 Experiments on Multi-Part Dataset
Implementation Details. On the DigitalTwinArt‑PM Dataset, we compare against PARIS* (PARIS augmented with depth supervision), DigitalTwinArt[65], and ArtGS[38]. On the GS‑PM Dataset, we compare against DigitalTwinArt[65] and ArtGS[38]. We report the mean for each metric over the 5 trials at the high-visibility state.
@@ 18 MPArt-90
Fig. 15 shows the statistics of our MPArt-90 benchmark. It contains 90 objects from 20 categories, including 12 objects from PARIS Two-Part Dataset [33] and 2 objects from DigitalTwinArt-PM Dataset [65]. Fig. 16 shows some examples of MPArt-90.

### v2

@@ 16.1 Experiments on PARIS Two-Part Dataset
Implementation Details. Following [weng2024digitaltwin], we compare against Ditto [Jiang_2022_ditto], PARIS [liu2023paris], PARIS* (PARIS augumented with depth supervision), CSG-reg [weng2024digitaltwin], 3Dseg-reg [weng2024digitaltwin], DigitalTwinArt [weng2024digitaltwin], and ArtGS [liu2025artgs]. For a fair evaluation, we follow DigitalTwinArt and report the mean std for each metric over the 10 trials at the high-visibility state.
@@ 16.2 Experiments on Multi-Part Dataset
Implementation Details. On the DigitalTwinArt‑PM Dataset, we compare against PARIS* (PARIS augmented with depth supervision), DigitalTwinArt[weng2024digitaltwin], and ArtGS[liu2025artgs]. On the GS‑PM Dataset, we compare against DigitalTwinArt[weng2024digitaltwin] and ArtGS[liu2025artgs]. We report the mean for each metric over the 5 trials at the high-visibility state.
@@ 18 MPArt-90
Tab. 7 shows the statistics of our MPArt-90 benchmark. It contains 90 objects from 20 categories, including 12 objects from PARIS Two-Part Dataset [liu2023paris] and 2 objects from DigitalTwinArt-PM Dataset [weng2024digitaltwin]. Fig. 15 shows some examples of MPArt-90.
| 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 15 | 20 | Overall | |
| Storage | 9 | 9 | 7 | 5 | – | 3 | – | 1 | – | 1 | 1 | 36 |
| Table | 5 | 7 | 5 | 4 | 1 | – | 1 | 2 | 1 | – | – | 26 |
| Blade | 2 | – | – | – | – | – | – | – | – | – | – | 3 |
| Fridge | 2 | 1 | – | – | – | – | – | – | – | – | – | 3 |
| Oven | 1 | 2 | – | – | – | – | – | – | – | – | – | 3 |
| Window | – | – | 1 | – | – | – | – | – | – | – | – | 1 |
| Washer | 2 | – | – | – | – | – | – | – | – | – | – | 2 |
| Foldchair | 2 | – | – | – | – | – | – | – | – | – | – | 2 |
| Door | – | 2 | – | – | – | – | – | – | – | – | – | 2 |
| Bucket | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Suitcase | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Box | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| DishWasher | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Safe | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Laptop | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Toilet | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| USB | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Scissor | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Stapler | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Microwave | 1 | – | – | – | – | – | – | – | – | – | – | 1 |
| Overall | 34 | 21 | 14 | 10 | 1 | 3 | 1 | 3 | 1 | 1 | 1 | 90 |